# UGG: Unified Generative Grasping

### v1

@@ 1 ${\bm{h}}_{T},{\bm{c}}_{T}\sim\mathcal{N}({\bm{0}},\mathbf{I})$ for $t$ = $T,\ldots,1$ do
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] is the concatenation.
A.2 Generation Algorithms
Appendix B Implementation
B.1 Structure of UGG
@@
B.2 Unified Shape Encoding
@@
Hands. As previously mentioned, we exclusively encode hand joint parameters ${\bm{\theta}}$ into the latent space using the dedicated hand VAE. The hand VAE comprises an encoder with a multilayer perceptron (MLP) of sizes [22, 32, 64] and a decoder with an MLP of sizes [64, 32, 22]. Both the encoder and decoder feature ReLU and BatchNorm between two linear layers. Each joint parameter is normalized to the range $[-1,1]$ before being inputted into the encoder, based on the minimum and maximum values in the data set.
For hand rotation ${\bm{R}}$, we adopt the flattened 6D rotation representation [82]. Transformation ${\bm{t}}$ is scaled from the range $[-0.3,0.3]$ to $[-1,1]$.
@@
B.3 Test-time Optimization
@@ B.5 Model Parameters
Appendix C Experiment Details
C.1 Dataset Setup
The DexGraspNet [71] dataset consists of 5355 objects, each associated with over 200 grasps based on 5 scales $\left\{0.06,0.08,0.1,0.12,0.15\right\}$. Our train/test splits adhere to those provided by the dataset. Although the original data set lacks specific information about the novel object category, we manually label the novel object category based on the provided object name code. During testing, we evenly distribute the objects of each scale.
@@ C.1 Dataset Setup
C.2 Training and Inference Time
The training of LION VAE consumes $\approx 560$ GPU hours, whereas VAE training at the hand takes $\approx 1.3$ GPU hours. The training of Physics Discriminator takes $\approx 10$ GPU hours.
@@ C.2 Training and Inference Time
For the UGG model on all DexGraspNet, the training time amounts to $\approx 4000$ GPU hours for $3000$ epochs. Training in the 20 objects subset requires $\approx 192$ GPU hours, and on the 10 bottles subset, it takes $\approx 96$ GPU hours. The inference time to generate $200$ grasps (equivalent to $200$ objects and $200$ pairs of grasp and object generation) is roughly $66$ seconds, and optimization for $100$ steps requires approximately $12$ seconds.
@@
Figure 13: Visualization of grasps for a single object at various scales. Each row features the same object with five different scales in the dataset, while each column corresponds to one scale.
Figure 14: Data distribution and grasp generation success rate (w/o discriminator) based on object scales.
Appendix D Additional Results
D.1 Grasp Generation of Different Scales
The DexGraspNet dataset comprises five scales for each object: $\left\{0.06,0.08,0.1,0.12,0.15\right\}$. Fig. 13 illustrates the grasp generation results for one object across these five scales. An intriguing observation emerges as the grasps on smaller-scale objects appear to be visually superior compared to those on the largest scale. To delve deeper, we analyze the dataset, examining the proportion of the data from each scale and the success rate of the grasps generated on each scale in Fig. 14. The dataset manifests an imbalance, with a higher representation of smaller-scale objects (0.06, 0.08, 0.1) and a lower representation of large-scale objects. The portion of 0.15-scale objects is merely half that of 0.08-scale objects. This skewed distribution potentially leads the model to emphasize learning from smaller-scale objects, a phenomenon reflected in the model’s performance in generating grasps for different object scales. The success rate of 0.15 scale objects is only half that of 0.06 scale objects. Furthermore, this impact extends to object generation results, as is evident in the model trained on the entire dataset, which tends to generate small, cube-like objects.

### v2

@@ 1 ${\bm{ohc}}_{T}=({\bm{o}}_{T},{\bm{h}}_{T},{\bm{c}}_{T})\sim\mathcal{N}({\bm{0}},\mathbf{I})$
5       ${\bm{\epsilon}}={\bm{\epsilon}}_{\xi}({\bm{ohc}}_{t},{\bm{t}})$
6       ${\bm{ohc}}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left({\bm{ohc}}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\bm{\epsilon}}\right)+\sigma_{t}{\bm{z}}$
7
@@ 0.A.2 Generation Algorithms
Appendix 0.B Implementation
0.B.1 Structure of UGG
@@
0.B.2 Unified Shape Encoding
@@
Hands. As previously mentioned, we exclusively encode hand joint parameters ${\bm{\theta}}$ into the latent space using the dedicated hand VAE. The hand VAE comprises an encoder with a multilayer perceptron (MLP) of sizes [22, 32, 64] and a decoder with an MLP of sizes [64, 32, 22]. Both the encoder and decoder feature ReLU and BatchNorm between two linear layers. Each joint parameter is normalized to the range $[-1,1]$ before being inputted into the encoder, based on the minimum and maximum values in the data set. For hand rotation ${\bm{R}}$, we adopt the flattened 6D rotation representation [87]. Transformation ${\bm{t}}$ is scaled from the range $[-0.3,0.3]$ to $[-1,1]$.
@@
0.B.3 Test-time Optimization
@@ 0.B.5 Model Parameters
Appendix 0.C Experiment Details
0.C.1 Dataset Setup
DexGraspNet. The DexGraspNet [76] dataset consists of 5355 objects, each associated with over 200 grasps based on 5 scales $\left\{0.06,0.08,0.1,0.12,0.15\right\}$. Our train/test splits adhere to those provided by the dataset. Although the original data set lacks specific information about the novel object category, we manually label the novel object category based on the provided object name code. During testing, we evenly distribute the objects of each scale.
@@ 0.C.1 Dataset Setup
HO3D. The HO3D dataset [19] stands as a comprehensive resource for human-object interaction research, including hand pose estimation, reconstruction, and generation. is a general human-object interaction dataset built for hand-object interaction estimation, reconstruction, and generation. It comprises 10 3D objects sourced from the YCB dataset. We include all 10 objects in our evaluation, following the procedure outlined in [24]. These 10 objects do not appear in the DexGraspNet dataset, and there may be overlaps in the categories they belong to.
GRAB. The GRAB dataset [69] features whole-body grasping actions performed by 10 subjects on 51 objects selected from the YCB dataset. Following prior works [25, 38], we use 6 test objects for evaluation. Similarly, these objects do not appear in the DexGraspNet dataset, and there may be overlaps in the categories they belong to.
0.C.2 Metrics
In our evaluation on the DexGraspNet, we adhere to the metrics established in their paper [76]. On HO3D and GRAB, we follow open sourced resources of [38, 25, 24], ensuring a fair comparison with baseline methods. Metrics are based on two key aspects, quality, and diversity, and we introduce each metric used in these benchmarks below.
Quality. 1) Success rate (%): A grasp is considered successful if it can endure at least one of the six gravity directions in the Isaac Gym simulator and exhibits a maximal penetration depth of less than 0.5 cm. 2) Mean $Q_{1}$: $Q_{1}$ [16] is defined as the radius of the inscribed sphere of ConvexHull ($\cup_{i}w_{i}$), representing the norm of the smallest wrench that can destabilize the grasp. The contact threshold is set to 1cm. 3) Maximal penetration depth (cm): The maximal penetration depth is calculated as the penetration depth from the object point cloud to the hand meshes. 4) Penetration Volume (cm3): The penetration volume between object and hand is computed by voxelizing the object mesh into 1 mm3 cubes and measuring overlapping voxels. 5) Simulation Displacement (cm): By putting both object and hand into the PyBullet [12] simulator, the displacement is measured by the displacement of object’s center of mass under the influence of the gravity.
Diversity. 1) Diversity on the DexGraspNet is evaluated through joint angle entropy. The joint motion range is discretized into 100 bins, and all generated samples contribute to estimating a probability distribution. The entropy is then computed as the entropy of these distributions. The reported value represents the mean and standard deviation across all joints. 2) Diversity on the HO3D and GRAB are computed by first clustering generated grasps into 20 clusters using K-means and measuring the entropy of cluster assignments.
Given that DexGraspNet relies on the ShadowHand [61] model (as does the UGG model), whereas the HO3D and GRAB datasets are built upon the MANO [57] hand model (so do the compared methods on these datasets), we briefly explore the influence of the hand model on these metrics. Since the compared methods on DexGraspNet are all implemented using the ShadowHand model, there is no need to address any influence there. The quality metrics on the HO3D and GRAB datasets are computed on both hand and object mesh. As both hand models offer mesh representations, their impact is minimal. However, the diversity metric of the MANO hand relies on hand keypoints, while that of the ShadowHand is based on hand joint angles. Notably, the former boasts over twice the number of parameters compared to the latter, thus giving an advantage to the MANO hand in terms of diversity metric.
0.C.3 Baseline Methods
We quote the results of the baseline methods on the HO3D and GRAB datasets from the ContactGen [38] paper, while the results of DDG and GraspTTA are taken from the DexGraspNet [76] paper.
Now we introduce how we conducted tests on UniDexGrasp [82]. UniDexGrasp encompasses two stages: a grasp generation stage and a reinforcement learning stage. For our assessment, we exclude the reinforcement learning part and focus only on evaluating the results of their grasp synthesis model. A notable difference in the task is that UniDexGrasp addresses how to grasp an object from the table, whereas other works primarily focus on how the generated pose can grasp (or more precisely, hold) the given object. Such distinction leads to two considerations in implementation and result interpretation. Firstly, the input of UniDexGrasp requires a table (positioned on the x-y plane with z=0), an extra input compared to all other methods. We opt to retain this difference to ensure a fair and transparent comparison by reusing the open-sourced model. We do not consider the table when calculating the metrics, therefore, the testing pipeline can mimic a similar environment as other methods. Secondly, the objective of UniDexGrasp is more about providing direction for the subsequent RL part to learn to grasp an object form the table. Therefore, it is acceptable to position slightly further away from the object. Such differences in the objective will result in lower penetration and simulation success rates, as reflected in the numbers reported.
0.C.4 Training and Inference Time
All experiments are conducted on a machine with 8 NVIDIA Tesla A100s. Running time listed below is based on NVIDIA Tesla A100 GPU.
@@ 0.C.4 Training and Inference Time
The training of LION VAE consumes $\approx 560$ GPU hours, whereas VAE training at the hand takes $\approx 1.3$ GPU hours. The training of Physics Discriminator takes $\approx 10$ GPU hours. For the UGG model on the entire DexGraspNet, the training time amounts to $\approx 4000$ GPU hours for $3000$ epochs. Training in the 20 objects subset requires $\approx 192$ GPU hours, and on the 10 bottles subset, it takes $\approx 96$ GPU hours. The inference time to generate $200$ grasps on one GPU (equivalent to $200$ objects and $200$ pairs of grasp and object generation) is roughly $66$ seconds, and optimization for $100$ steps requires approximately $12$ seconds.
Appendix 0.D Additional Results
Table 4: Quantitative results of grasp generation of UGG and benchmark methods on the HO3D and GRAB datasets.
HO3D GRAB
Penetration Simulation Entropy Penetration Simulation Entropy
[…]
GraspTTA [24] 7.37 5.34 2.70 - - -
GrabNet [69] 15.50 2.34 2.80 3.65 1.72 2.72
GF [26] 93.01 - 2.75 - - -
HALO [25] 25.84 3.02 2.81 3.61 2.09 2.88
ContactGen [38] 9.96 2.70 2.81 2.72 2.16 2.88
UGG (ours) 1.23 2.11 2.98 1.08 1.37 2.98
0.D.1 Grasp Generation on Human-object Interaction Datasets
We test our method on the human-object interaction datasets HO3D [19] and GRAB [69, 5], which contains 10 and 6 test objects respectively. Given the misalignment in the hand model, we choose not to finetune our model on these datasets but rather test its generalization ability. Comparisons on the HO3D and GRAB datasets involves GrabNet [69], HALO [25], GraspTTA [24], GraspingField [26], and ContactGen [38]. Despite these datasets utilizing a different hand model from the one employed in the DexGraspNet dataset, we argue that the applied metrics remain comparable (See Sec. 0.C.2).
As presented in Tab. 4, UGG achieves notably lower penetration volume and simulation displacement, and higher entropy in comparison to the prior methods. Moreover, such performances are achieved without finetuning, and the majority of the objects in these datasets are absent from the DexGraspNet training set.
Table 5: Success rate breakdown of GraspTTA, UDG, and UGG (ours) on DexGraspNet.
GraspTTA UDG UGG
generation only 4.2 10.6 43.6
+TTA 24.5 23.3 64.1
+Physics Discriminator 28.6 26.9 72.7
0.D.2 Detailed Break-down on Model Performance
We present a detailed performance breakdown of how each module of the compared models - GraspTTA, UniDexGrasp, and UGG - performs on DexGraspNet. We also implemented the Physics Discriminator for two compared methods. Since each model includes a generation module, a test-time adaptation module, and the Physics Discriminator, we evaluate their performance by sequentially adding each module.
The results, detailed in Table 5, show that our method achieves the best performance among all generative models. The addition of our proposed physics discriminator consistently improves performance across all variants. However, it is noted that a weaker base model will naturally result in lower overall success rates, as we must select 75% of the grasps, and the top grasps for these models are not as strong. Additionally, models like GraspTTA and UGG, which include contact information, benefit more from test-time adaptation due to the informative nature of the contact data.
@@ 0.D.2 Detailed Break-down on Model Performance
Figure 13: Grasp Generation of different scales. (a) Visualization of grasps for a single object at various scales. Each row features the same object with five different scales in the dataset, while each column corresponds to one scale. (b) Data distribution and grasp generation success rate (w/o discriminator) based on object scales.
0.D.3 Grasp Generation of Different Scales
[…]
0.D.4 Qualitative Comparison
Figure 14: Visualization of qualitative comparison of grasps generated by three models on the DexGraspNet objects, UGG (top), UniDexGrasp (mid), and GraspTTA (bottom).
Fig. 14 shows comparison of qualitative results of three methods. While most grasps are considered successful as per evaluation metric, UGG’s grasps are visually exihibiting more precise contact control. Grasps generated by UniDexGrasp are much further away from the object, which aligns with our observation of the goal of the UniDexGrasp explained in Sec. 0.C.3. Some examples of GraspTTA suffers from unavoidable penetration, which attribute to the limited initial generation result from CVAE.
Appendix 0.E Potential Impact
Given the substantial time and resource investment required for training this model, code and model will be made publicly available. We find no clear negative social impact associated with this work; however, we acknowledge the necessity of handling the generative model with care to prevent any potential harm during execution. We believe that the challenge addressed in this paper, the generation of objects based on hand gestures, presents new possibilities for human-centric object design that could greatly benefit society as a whole.